OpenStudy (rayford):

What are the solution(s) to the following system of equations? y = x2 – 1 y = x – 1

OpenStudy (lgbasallote):

try subtracting the equations...what do you get?

OpenStudy (anonymous):

\[x ^{2}-1=x-1\] \[x ^{2}-x=0\] \[x(x-1)=0\]

OpenStudy (rayford):

so whats the answer?